### v3

Abstract
Most existing methods rely on the similarity correlation matrix to characterize the repetitiveness of actions, but their scalability is hindered due to the quadratic computational complexity. In this work, we introduce a novel approach that employs an action query representation to localize repeated action cycles with linear computational complexity. Based on this representation, we develop two key components to tackle the essential challenges of temporal repetition counting. Firstly, to facilitate open-set action counting, we propose the dynamic update scheme on action queries. Unlike static action queries, this approach dynamically embeds video features into action queries, offering a more flexible and generalizable representation. Secondly, to distinguish between actions of interest and background noise actions, we incorporate inter-query contrastive learning to regularize the video representations corresponding to different action queries. As a result, our method significantly outperforms previous works, particularly in terms of long video sequences, unseen actions, and actions at various speeds. On the challenging RepCountA benchmark, we outperform the state-of-the-art method TransRAC by 26.5% in OBO accuracy, with a 22.7% mean error decrease and 94.1% computational burden reduction. Code will be publicly available.
keywords:
Temporal repetition counting, Video understanding
1 Introduction
Temporal periodicity is a ubiquitous phenomenon in the natural world. Temporal Repetition Counting (TRC) aims to accurately measure the number of repetitive action cycles within a given video and holds significant potential for applications such as fitness monitoring [11] and motion generation [49].
Figure 1: Conceptual workflow comparison of (a) similarity-matrix-based methods and (b) our proposed action query-based method. (a) Most existing methods use similarity matrices calculated between each frame to detect repetitive actions, resulting in a time complexity of $O(T^{2}C)$, where $T$ denotes video length and $C$ denotes feature dimension. (b) On the other hand, our method employs action queries to represent each action cycle and estimates their classes and temporal locations. Each query is classified into either “repetitive actions” or “others ($\emptyset$)” class labels. The temporal location is expressed by the midpoint and duration on the timeline. During training, we perform bipartite matching to uniquely associate a prediction to a GT action cycle, taking into account both class labels and temporal locations. Predictions with no match should yield a $\emptyset$ class prediction, indicating that these are not repetitive actions. This novel formulation reduces the complexity from quadratic to linear as $O(TC)$.
Pioneer methods  [17, 1, 7, 39, 29, 38, 6] represent time-series video data as one-dimensional signals and employ spectral analysis techniques such as the Fourier transform. While suitable for short videos with fixed periodic cycle lengths, these methods struggle to handle real-world scenarios with varying cycle lengths and sudden interruptions. Recent studies shift to deep learning-based methods [18, 9, 14, 47, 19] and show promising performances. Notably, most of these methods, such as RepNet [9] and TransRAC [14], utilize a temporal similarity correlation matrix to depict repetitiveness, as illustrated in Fig. 1 (a). Nevertheless, the computational complexity of this representation grows quadratically with the number of input frames $T$, highlighting a significant gap in scalability that hinders their application to real-world scenarios of varying action periods and dynamics.
Recent progress in action detection [34, 24] introduces an efficient representation of action periods by associating each action instance with an action query, similar to DETR [3]. Inspired by this, we propose to formulate the TRC problem as a set prediction task where the goal is to detect every action cycle by representing it as an action query, as illustrated in Fig. 1 (b). Based on this query representation, we use a Transformer encoder-decoder network [50] to detect repetitive action instances and their temporal positions, defined by their midpoints and durations. This novel formulation reduces the complexity from quadratic to linear  [^1] In implementation, we employ deformable attention modules as proposed in DeformableDETR [50]. and enables counting long videos with varying action periods. However, directly applying the action detection approaches [24, 45, 34] to the TRC problem proves inadequate (Tab. 1) in addressing two distinctive challenges unique to TRC. These challenges underscore the complexity of TRC, highlighting why TRC is not merely another action detection task but requires a nuanced approach that considers the unique nature of repetitive actions. We highlight the two inherent differences between TRC and the classical action detection task:
  1. 1.
    TRC requires recognizing open-set action instances depending on the input video, rather than detecting predefined action classes in the detection task.
  2. 2.
    TRC requires recognizing action instances with identical content, while detection does not.
As a result, approaching TRC as a simple action detection task results in inferior performance as shown in Sec. 4.3. In contrast, we propose two novel strategies to address these challenges and redefine the framework for TRC. In response to the first challenge, we propose Dynamic Action Query (DAQ), which adaptively updates the action query using content features extracted from the video (Fig. 2). This mechanism allows the decoder to attend to the repetitive actions based on the input video contents in a dynamic and contextually aware manner, without the need for manual definition. To tackle the second challenge, we define two action classes: “repetitive actions”, which are dynamically defined by DAQ, and the “others ($\emptyset$)” class. We further propose Inter-query Contrastive Learning (ICL) which ensures that primary repetitive action cycles are grouped together in the learned representation space while being separated from other distractors, such as background noise. The integration of two core components (DAQ and ICL) ensures that the action instances are identified adaptively based on the video content and their contextual similarity. In other words, our queries are designed to localize the contextually similar action instances, which aligns exactly with the definition of repetition counting. Extensive experiments validate the effectiveness of the two proposed designs.
We summarize our contributions as follows:
  1. 1.
    We provide a novel perspective to tackle the TRC problem using a simple yet effective representation for action cycles. Our approach reduces the computational complexity from quadratic to linear and is robust to varying action periods and video lengths.
  2. 2.
    We propose Dynamic Action Query to guide the model to focus on the repetitive actions contextually defined by the video content thereby improving generalization ability across different actions.
  3. 3.
    We introduce Inter-query Contrastive Learning to facilitate learning primary repetitive action representations and to distinguish them from distractions.
  4. 4.
    By addressing the scalability issue and unique challenges of TRC, our method on two challenging benchmarks notably surpasses state-of-the-art (SOTA) methods in terms of both accuracy and efficiency. Notably, our method strikes an effective balance in handling various action periods and video lengths, offering a significant leap forward in the practical application of TRC technologies.
2 Related Work
2.1 Temporal Repetition Counting
Traditional methods [17, 1, 7, 39, 29, 38, 6] frequently employ spectral or frequency domain techniques for the analysis of repetitive sequences, thereby preserving the underlying repetitive motion structures. While these conventional approaches are capable of effectively handling simple motion sequences or those characterized by fixed periodicity, they prove inadequate when confronted with non-stationary motion sequences encountered in real-world scenarios. In contrast, deep-learning-based approaches [18, 9, 14, 47, 19] have demonstrated remarkable performance improvements. Notably, RepNet [9] and TransRAC [14] leverage temporal similarity matrices of actions to construct models for counting temporal repetitions. However, these similarity-matrix-based methods are not scalable for long videos due to their quadratic computational complexity. Another research line involves predicting the start and end points of each cycle [47] from coarse to fine. Nevertheless, its practicality is hindered by the requirement for over 30 forward passes to count iteratively from a single video. In this paper, we introduce an effective action cycle representation by leveraging a Transformer encoder-decoder, which reduces the computational complexity from quadratic to linear and demonstrates superior performance in handling both fast and slow actions.
2.2 Temporal Action Detection
The field of temporal action detection [31, 48, 5, 45, 22] is typically classified into two categories: anchor-based methods, and anchor-free methods. Anchor-based methods [44, 20, 30] generate multiple anchors, subsequently classifying these anchors to determine the action boundaries. Anchor-free methods  [2, 35, 43, 21] predict action instances by directly regressing the boundary and the center point of an action instance. With the rapid development of Transformer technology, DETR [3] is introduced for object detection task [50, 28, 23, 46] and gains increasing popularity with promising performance. This paradigm promotes the study in many fields such as the action detection tasks [24, 37, 40, 42]. These methods establish a direct connection between action queries and the predicted action instances, enabling them to accurately predict the temporal boundaries of actions. Inspired by these promising results, we explore the possibility of utilizing a novel action query to represent the action cycle in TRC task. In contrast to existing action detection methods, our approach allows the model to capture the inherent repetitive content of an action cycle without relying on predefined class labels and effectively addresses confounding factors such as non-repetitive video backgrounds. This makes our approach well-suited for tackling the challenges of the TRC problem.
3 Method
3.1 Preliminary
DETR [3] is a pioneering object detection framework that builds upon the Transformer encoder-decoder architecture [40]. The overall DETR architecture [3] consists of three main components: a backbone to extract image features, an encoder-decoder Transformer, and the detection heads, i.e. feed-forward network (FFN) that makes the final detection prediction. The main features of DETR are the conjunction of a bipartite matching loss and transformers with (non-autoregressive) parallel decoding. The bipartite matching loss is a set-based Hungarian loss during training, which uniquely assigns a prediction to a GT object, and is invariant to a permutation of predicted objects. This design enables DETR to perform parallel processing and predict all objects simultaneously. We briefly review the workflow as follows.
Given the input image feature maps extracted by a CNN backbone, e.g. ResNet [13], DETR exploits a standard Transformer encoder-decoder network to transform the feature maps to be features of a set of object queries. An FFN and a linear projection are added on top of the object query features as the detection heads. The FFN acts as the regression branch that predicts the bounding box coordinates, i.e. box center coordinates, box height and width. The linear projection acts as the classification branch to produce the classification results, i.e. $\mathtt{object}$ vs. $\mathtt{no\,object}\,(\emptyset)$. The $\emptyset$ class is used to represent that no object is detected, playing a similar role to the “background” class in the standard object detection approaches.
DETR infers a fixed-size set of $N$ predictions in a single pass, where $N$ is set to be significantly larger than the typical number of objects in an image. During training, a Hungarian loss produces an optimal bipartite matching between predicted and GT objects and then optimizes the object position-specific losses. The matching procedure takes into account both the class prediction and the similarity of predicted and GT boxes and finds one-to-one matching for direct set prediction without duplicates.
Figure 2: Overview and detailed architecture design of our method. (Left) We utilize a DETR-inspired framework to predict the number of repetitive action cycles in an RGB video sequence. The architecture consists of a backbone network $\Phi(\cdot)$ for feature extraction, an encoder-decoder Transformer for processing query features, and prediction heads for action classification. By counting the queries classified as “repetitive actions”, we get the total count $N$. (Right) The detailed workflow within the black box on the left. Given the video features $\mathbf{F}$ as input, the encoder $\mathcal{E}(\cdot)$ produces two sets of query features, i.e. action queries and position queries. A query selection module then screens these queries, retaining only the most relevant ones for forwarding to the decoder $\mathcal{D}(\cdot)$. This process utilizes action classification results from prediction heads positioned after the encoder. The decoder adopts the Dynamic Action Query (DAQ) strategy, utilizing the selected action queries as priors, while incorporating a learnable position query and attending to the encoder’s output. The output embeddings from the decoder are then passed to their corresponding prediction head, which estimates the class label $\mathbf{A}$ and temporal location $\mathbf{P}$. During training, we employ bipartite matching to uniquely pair each prediction with a GT action cycle $\hat{\mathbf{P}}$, considering both class labels and temporal locations. Predictions that fail to match should yield $\emptyset$ class, signifying that they do not represent repetitive actions. The DAQ strategy addresses the open-set problem by contextually defining “repetitive actions” based on video features. To further distinguish these actions from $\emptyset$ distractions, we propose Inter-query Contrastive Learning (ICL). By integrating DAQ and ICL, our method effectively identifies contextually similar actions and excludes distractions, making it well-suited for tackling the challenges of the TRC problem.
3.2 Overview
Given an RGB video sequence with $T$ frames, the TRC task aims to predict an integer $N$ indicating the number of detected repetitive action cycles. Drawing inspiration from DETR [3], we streamline the problem as a set detection task in the temporal domain. The overall framework consists of three main components: a backbone network $\Phi(\cdot)$, an encoder-decoder Transformer, and prediction heads, as shown in Fig. 2. The backbone extracts video features $\mathbf{F}$, which is transformed to features of a set of action queries by an encoder $\mathcal{E}(\cdot)$ and a decoder $\mathcal{D}(\cdot)$. Two prediction heads are added on top of the query features. The prediction heads consist of an action head and a position head. The action head classifies the query, while the position head estimates its temporal location.
Building on previous enhancements in DETR models [23], we introduce two distinct types of queries to enhance the query-to-feature similarity and accelerate convergence: action queries and position queries. Action queries are designed to capture features pertinent to action classification, whereas position queries concentrate on the temporal dimensions of an action. As depicted in the detailed diagram on the right side of Fig. 2, two specialized prediction heads (i.e. action head and position head) are tasked with processing these respective sets of queries to generate predictions. To further accelerate convergence, we introduce a post-encoder query selection module that acts as an initial sieve, filtering out the majority of queries unlikely to represent repetitive actions before forwarding the most promising ones to the decoder. This enhances the efficiency and focus of the subsequent prediction process.
Considering the two challenges of the TRC tasks discussed in Sec. 1, we propose Dynamic Action Query (DAQ) strategy to address the open-set problem. DAQ directly uses dynamically updated action query features from the encoder to initialize the action query of the decoder. This strategy defines what “repetitive actions” should be detected based on the video content features in a dynamic and contextually aware manner. To tackle the second challenge, we define two action classes: “repetitive actions” and the “others ($\emptyset$)” class. We further propose Inter-query Contrastive Learning (ICL) to cluster repetitive action queries into a positive action set ($S^{+}$) and group the other queries to a negative set ($S^{-}$) in the feature space. Integrating DAQ and ICL allows our method to identify contextually similar action instances that are adaptively based on the video content, and exclude other distracting actions at the same time.
Similar to DETR [3], during training, we employ bipartite matching which uniquely assigns a prediction to a GT action cycle. The matching procedure takes into account both the class prediction and the similarity of predicted and GT temporal locations. We then optimize the action cycle-specific losses. In inference, our method produces a fixed-size set of $Q$ predictions in a single pass, where $Q$ is set to be significantly larger than the typical number of action cycles in a $T$-frame video sequence. By counting the queries classified as “repetitive actions”, we get the final total count value $N$.
In the following, we will introduce the network architecture design in Sec. 3.3, the DAQ and ICL modules in Sec. 3.4 and Sec. 3.5, and the model training in Sec. 3.6.
3.3 Model Architecture
Backbone. The backbone network $\Phi(\cdot)$ takes a sequence of $T$ video frames as input and extracts feature vectors $\mathbf{F}\in\mathbb{R}^{T\times C}$ for each frame, where $C$ denotes the feature dimension.
Encoder. The encoder $\mathcal{E}(\cdot)$ is a classical Transformer [40] architecture which has ${L}^{\text{enc}}$ standard encoder layers. The encoder transforms the video features $\mathbf{F}$ into two distinct query types: action queries $\tilde{\mathbf{E}}^{\text{act}}\in\mathbb{R}^{T\times C}$ and position queries $\tilde{\mathbf{E}}^{\text{pos}}\in\mathbb{R}^{T\times C}$, as shown in Fig. 2 (right). This design enhances the query-to-feature similarity and accelerates convergence, where action queries capture features pertinent to action classification. On the other hand, position queries concentrate on the temporal dimensions of an action. We defer to the supplementary material the detailed encoder architecture.
Query selection. We add a query selection module before passing the encoder output query features to the decoder. We route the query features outputted by the encoder to two respective prediction heads (described later in this section), which independently decode them into predictions for action class and temporal location, as depicted on the right side of Fig. 2. The action head processes the action queries $\tilde{\mathbf{E}}^{\text{act}}$ to estimate the action class $\tilde{\mathbf{A}}$ for each query, while the position head decodes the position queries $\tilde{\mathbf{E}}^{\text{pos}}$ to their temporal locations $\tilde{\mathbf{P}}$. Based on the prediction results, we discard query features classified as “others ($\emptyset$)”, retaining only those identified as “repetitive actions”. For the remaining queries, we rank them based on their classification confidence and preserve only the top $Q$ high-confidence queries. The others are discarded. After this selection process, both the action query features $\mathbf{E}^{\text{act}}$ and position query features $\mathbf{E}^{\text{pos}}$ have a dimension of $Q\times C$. Additional details will be provided when we discuss the prediction head later in this section.
Decoder. The decoder $\mathcal{D}(\cdot)$ is also a classical Transformer [40] architecture, consisting of ${L}^{\text{dec}}$ standard decoder layers. It processes a set of action queries $\mathbf{Q}^{\text{act}}\in\mathbb{R}^{Q\times C}$ and position queries $\mathbf{Q}^{\text{pos}}\in\mathbb{R}^{Q\times C}$, while simultaneously attending to the direct output (i.e. $\tilde{\mathbf{E}}^{\text{act}}$ and $\tilde{\mathbf{E}}^{\text{pos}}$) from the encoder, as depicted in Fig. 2 (right). Following the practice in DETR [3], these inputs are decoded in parallel across each decoder layer and transformed into the corresponding query features $\mathbf{D}^{\text{act}}\in\mathbb{R}^{Q\times C}$ and $\mathbf{D}^{\text{pos}}\in\mathbb{R}^{Q\times C}$. The input position queries $\mathbf{Q}^{\text{pos}}$ are initialized as learnable parameters. To address the open-set challenge in the TRC task, we introduce the DAQ strategy, which initializes the decoder input action queries using the selected encoder embeddings $\mathbf{E}^{\text{act}}$, i.e. $\mathbf{Q}^{\text{act}}=\mathbf{E}^{\text{act}}$. The details of the DAQ strategy will be elaborated in Sec. 3.4. Using self- and encoder-decoder attention over these embeddings, our method globally reasons selected video frame features together while being able to use the whole frame-wise video feature as context. The two sets of $Q$ queries are then independently transformed into corresponding prediction results, i.e. an action head decodes $\mathbf{D}^{\text{act}}$ into the action class prediction $\mathbf{A}$, and a position head decodes $\mathbf{D}^{\text{pos}}$ into the temporal location prediction $\mathbf{P}$. We detail the two heads in the following subsection.
Prediction heads. The prediction heads are designed to make the final predictions using the features processed by the encoder-decoder. They comprise two networks: an action head and a position head, both of which are Multi-Layer Perceptrons (MLPs).
The action head processes the output action queries from the decoder $\mathbf{D}^{\text{act}}$ to estimate the action class $\mathbf{A}\in\mathbb{R}^{Q\times 1}$ for each query using a softmax function. Each element of $\mathbf{A}$ represents the probability that the corresponding query is a “repetitive action”, with values spanning from 0 to 1. In practice, a threshold value, $\alpha$, is used to categorize the queries: values greater than $\alpha$ classify a query as a “repetitive action” while values less than or equal to $\alpha$ categorize it as “others ($\emptyset$)”. Recall that in the query selection module, we rank the queries based on their estimated probabilities $\tilde{\mathbf{A}}\in\mathbb{R}^{T\times 1}$, which serve as confidence scores. We then select the top $Q$ queries with the highest confidence scores according to the ranking.
The position head estimates the temporal location $\mathbf{P}=(\mathbf{m},\mathbf{d})\in\mathbb{R}^{Q\times 2}$ for each query based on the output position queries from the decoder $\mathbf{D}^{\text{pos}}$. We use the midpoint $\mathbf{m}\in\mathbb{R}^{Q}$ and the duration $\mathbf{d}\in\mathbb{R}^{Q}$ to denote the temporal location of an action cycle.
Bipartite matching. Following the practice of DETR [3], our method produces a fixed-size set of $Q$ predictions in a single pass, where $Q$ is set to be much larger than the typical number of action cycles in a $T$-frame video sequence. Note that the $Q$ predictions are in a random order. One of the main difficulties of training is to match the predicted action cycles (classes and temporal locations) to the GT action cycles, therefore, optimizing action cycle-specific losses, e.g. classification loss, temporal position loss, etc.
Let us denote GT repetitive action cycle labels by $\hat{\mathbf{Y}}=(\hat{\mathbf{A}},\hat{\mathbf{P}})$. $\hat{\mathbf{A}}=\mathbf{1}\in\mathbb{R}^{\hat{N}\times 1}$ denotes the $\hat{N}$ repetitive action class labels, where we assign the value 1 to the “repetitive actions”. $\hat{\mathbf{P}}\in\mathbb{R}^{\hat{N}\times 2}$ denotes the GT temporal positions of the $\hat{N}$ repetitive action, represented by $\hat{N}$ sets of midpoints and duration. Similarly, we denote the predicted repetitive actions by $\mathbf{Y}=(\mathbf{A},\mathbf{P})$, a set of size $Q$ predictions. We set $Q$ much larger than $\hat{N}$ empirically. We consider $\hat{\mathbf{Y}}$ also a set of size $Q$ padded with “others ($\emptyset$)”, i.e. we assign the value 0 to denote $\emptyset$ class. To find a bipartite matching between these two sets $\hat{\mathbf{Y}}$ and $\mathbf{Y}$, we search for a permutation of $Q$ elements $\sigma\in\mathfrak{S}_{Q}$ with the lowest cost:
$\hat{\sigma}=\arg\min_{\sigma\in\mathfrak{S}_{Q}}\sum_{i=1}^{Q}\mathcal{L}_{\text{match}}(\hat{\mathbf{Y}}_{i},\mathbf{Y}_{\sigma(i)}),$ (1)
where $\mathcal{L}_{\text{match}}$ is a pair-wise matching cost between GT action cycle $\hat{\mathbf{Y}}_{i}$ and a prediction cycle with index $\sigma(i)$.
The matching cost takes into account both the action classification result and the similarity of the predicted temporal locations and the GT temporal location. Each element $i$ of the GT action cycle set can be seen as $\hat{\mathbf{Y}}_{i}=(\hat{\mathbf{A}}_{i},\hat{\mathbf{P}}_{i})$ where $\hat{\mathbf{A}}_{i}$ is the target class label (1 for “repetitive actions”, and 0 for $\emptyset$) and $\hat{\mathbf{P}}_{i}\in\mathbb{R}^{2}$ is a vector that denotes the midpoint time position and lasting duration of a GT action cycle. For the prediction with index $\sigma(i)$, we define probability of class $\hat{\mathbf{A}}_{i}$ as $p_{\sigma(i)}(\hat{\mathbf{A}}_{i})$. Then we can define the matching cost as
$\begin{split}\mathcal{L}_{\text{match}}(\hat{\mathbf{Y}}_{i},\mathbf{Y}_{\sigma(i)})&=-\mathbbm{1}_{\{\hat{\mathbf{A}}_{i}\neq\emptyset\}}p_{\sigma(i)}(\hat{\mathbf{A}}_{i})\\ &+\mathbbm{1}_{\{\hat{\mathbf{A}}_{i}\neq\emptyset\}}\mathcal{L}_{\text{pos}}(\hat{\mathbf{P}}_{i},\mathbf{P}_{\sigma(i)}),\end{split}$ (2)
where $\mathbbm{1}$ is an indicator function.
To measure the similarity of the predicted temporal locations and the GT temporal location, we define $\mathcal{L}_{\text{pos}}$ using the linear combination of the $L_{1}$ distance and the generalized Intersection over Union (IoU) loss [32]. Overall, the position loss is defined as
$\begin{split}\mathcal{L}_{\text{pos}}(\hat{\mathbf{P}}_{i},\mathbf{P}_{\sigma(i)})&=\lambda_{\text{L1}}\|\hat{\mathbf{P}}_{i}-\mathbf{P}_{\sigma(i)}\|_{1}\\ &+\lambda_{\text{gIoU}}\mathcal{L}_{\text{gIoU}}(\hat{\mathbf{P}}_{i},\mathbf{P}_{\sigma(i)}),\end{split}$ (3)
where $\lambda_{\text{gIoU}},\lambda_{\text{L1}}\in\mathbb{R}$ are hyperparamters.
By employing the Hungarian matching algorithm [16] to optimize Eq. 1, we can achieve the final optimal matching $\hat{\sigma}$ which uniquely assigns a prediction to a GT action cycle, finding one-to-one matching without duplicates. Notice that the matching cost between a repetitive action instance and $\emptyset$ doesn’t depend on the prediction, which means that in that case the cost is a constant. Following this step, we can apply the corresponding losses related to classification and temporal location prediction.
We define a Hungarian loss for all pairs matched in the previous step to supervise both classification and temporal location predictions, i.e. a linear combination of a negative log-likelihood loss for class prediction and a temporal position loss defined in Eq. 3:
$\begin{split}\mathcal{L}_{\text{Hungarian}}(\hat{\mathbf{Y}},\mathbf{Y})&=\sum\limits_{i=1}^{Q}\left[-\text{log}\,{p_{\hat{\sigma}(i)}}{(\hat{\mathbf{A}}_{i})}\right.\\ &+\left.\mathbbm{1}_{\{\hat{\mathbf{A}}_{i}\neq\emptyset\}}\mathcal{L}_{\text{pos}}(\hat{\mathbf{P}}_{i},\mathbf{P}_{\hat{\sigma}(i)})\right],\end{split}$ (4)
where $\hat{\sigma}$ is the optimal assignment computed in the previous bipartite matching step, i.e. Eq. 1.
3.4 Dynamic Action Query
As discussed in Sec. 1, the TRC problem requires recognizing open-set action instances depending on the video content, where the action category is not predefined. Therefore, we propose the Dynamic Action Query strategy, which adaptively updates the action query $\mathbf{Q}^{\text{act}}$ using the selected encoder action query features $\mathbf{E}^{\text{act}}$, i.e. $\mathbf{Q}^{\text{act}}=\mathbf{E}^{\text{act}}$, as shown in Fig. 2. The DAQ strategy, simple yet effective, provides priors for the decoder to dynamically and contextually attend to “repetitive actions” based on the input video content, eliminating the need for manually defining action categories and thereby enhancing the model’s generalization capability. This mechanism ensures that the cross-attention module in the decoder pools features from the entire video while focusing on relevant local repetitive features. Such dynamic adjustment allows the model to concentrate on pertinent “repetitive actions”, offering a deeper understanding of action queries and improving training convergence.
In summary, the DAQ approach not only circumvents the need for manual action category definitions but also leverages dynamic, contextually aware priors to enhance both the focus and generalization capabilities of the model. Our supplementary material further explores various methods for initializing the queries in the decoder $\mathcal{D}(\cdot)$, confirming that the DAQ strategy is the most effective.
3.5 Inter-query Contrastive Learning
Since the input video may contain other distractors such as the background motion  [^2] We assume that each video contains only a primary repetitive action type, which is also the case for the public dataset., another unique challenge to the TRC task is to recognize action instances with identical content. This requires that the actions of interest we classify exhibit similarity in their motion patterns, while other action queries should have dissimilar representations. To tackle this challenge, we propose Inter-query Contrastive Learning to distinguish the action queries. Intuitively, based on the classification predictions $\mathbf{A}$, we partition the action queries decoded by the decoder $\mathbf{D}^{\text{act}}$ into two categories and employ inter-query contrastive learning on them, as shown in Fig. 2 (right). The features that are classified as “repetitive actions” form the positive set $S^{+}$, while the other features form the negative set $S^{-}$. Then we apply contrastive learning using InfoNCE loss [12] $\mathcal{L}_{\text{ctrs}}$ over the representation space:
$\begin{split}\mathcal{L}_{\text{ctrs}}&=-\sum\limits_{i\in S^{+}}{\rm log}\left(\frac{\mathcal{L}_{+}}{\mathcal{L}_{+}+\mathcal{L}_{-}}\right),\\ \mathcal{L}_{+}&=\sum\limits_{s\in S^{+},s\neq i}{\rm exp}(\mathbf{D}^{\text{act}}_{i}\cdot\mathbf{D}^{\text{act}}_{s})/\tau,\\ \mathcal{L}_{-}&=\sum\limits_{s\in S^{-}}{\rm exp}(\mathbf{D}^{\text{act}}_{i}\cdot\mathbf{D}^{\text{act}}_{s})/\tau,\end{split}$ (5)
where $\tau$ is the temperature parameter, and $\cdot$ denotes inner product.
3.6 Training
We train our model in an end-to-end manner using the overall loss function:
$\mathcal{L}=\lambda_{\text{Hungarian}}\mathcal{L}_{\text{Hungarian}}+\lambda_{\text{ctrs}}\mathcal{L}_{\text{ctrs}},$ (6)
where $\lambda_{\text{Hungarian}},\lambda_{\text{ctrs}}\in\mathbb{R}$ are the coefficients. Following DETR [3], we also found it helpful to use auxiliary losses in the decoder during training. Specifically, we add $\mathcal{L}_{\text{Hungarian}}$ loss on the predictions from the prediction heads right after the encoder i.e. $\tilde{\mathbf{A}}$ and $\tilde{\mathbf{P}}$, and add prediction heads and $\mathcal{L}_{\text{Hungarian}}$ loss after each decoder layer. All prediction heads share their parameters. This ensures the model focuses on the correct features at each stage consistently, thereby accelerating convergence.
4 Experiments
4.1 Datasets and Metrics
RepCountA dataset [14] is currently the largest and most challenging benchmark for the video TRC task  [^3] The RepCountB [14] test subset is proprietary and not publicly available.. It is primarily compiled from fitness videos on YouTube, including a wide range of fitness activities conducted in diverse settings, including homes, gyms, and outdoor environments. This dataset stands out due to its extensive video lengths, significant variations in the average motion cycle, and a higher number of repetitive cycles compared to prior datasets [9, 33, 18, 47]. We use the start and end positions of each action instance to compute the GT label $\hat{\mathbf{P}}$ provided by the annotations. We train our model on the RepCountA train set and select the best model on the validation set. We report the evaluation results on the test set.
Table 1: Comparison to the state-of-the-arts on RepCountA [14] dataset. We compare with SOTA action recognition/segmentation methods (top block), TRC methods (second block), and action detection methods (third block with $\dagger$). We further report MAE and OBO metrics for short-, medium-, and long-period test actions.
Backbone ${\rm MAE}\downarrow$ ${\rm OBO}\uparrow$ ${\rm MAE_{s}}\downarrow$ ${\rm OBO_{s}}\uparrow$ ${\rm MAE_{m}}\downarrow$ ${\rm OBO_{m}}\uparrow$ ${\rm MAE_{l}}\downarrow$ ${\rm OBO_{l}}\uparrow$
X3D [10] X3D 0.9105 0.1059 - - - - - -
TANet [26] TANet 0.6624 0.0993 - - - - - -
VideoSwinT [25] ViT 0.5756 0.1324 - - - - - -
GTRM [15] I3D 0.5267 0.1589 - - - - - -
RepNet [9] ResNet 0.5865 0.2450 0.7793 0.0930 0.5893 0.1591 0.4549 0.4062
Zhang et al. [47] 3D-ResNext 0.8786 0.1554 - - - - - -
TransRAC [14] ViT 0.4891 0.2781 0.5789 0.0233 0.4696 0.2955 0.4420 0.4375
Li et al. [19] ViT 0.3841 0.3860 - - - - - -
TadTR [24] $\dagger$ I3D 1.1314 0.0662 0.8364 0.0233 1.1591 0.0000 1.3106 0.1406
ActionFormer [45] $\dagger$ I3D 0.4990 0.2781 0.4164 0.1628 0.3768 0.3409 0.6385 0.3125
ReAct [34] $\dagger$ TSN 0.4592 0.3509 0.2805 0.1576 0.3037 0.4318 0.6862 0.3906
Ours TSN 0.2809 0.4570 0.2411 0.1628 0.1792 0.5455 0.3776 0.5938
Ours I3D 0.3305 0.4437 0.2421 0.2093 0.2012 0.5227 0.4788 0.5469
Ours ViT 0.2622 0.5430 0.2257 0.2558 0.2002 0.5909 0.3294 0.7031
UCFRep dataset [47] is a subset of the UCF101 dataset [36], including fitness videos and daily life videos. Following previous work [14, 19], we do not use the train set but directly test our model on the test set to evaluate the model generalization ability.
Metrics. Following previous works [14, 9, 47, 19], we compute two commonly used metrics, OBO and MAE, to evaluate the model performance. OBO (Off-By-One count error) measures the percentage that the predicted count is within the GT count ±$\pm 1$ range. MAE (Mean Absolute Error) measures the normalized absolute difference between the predicted and GT counts. Formally,
${\rm OBO}$ $=\frac{1}{M}\sum_{m=1}^{M}\left|N_{m}-\hat{N}_{m}\leq 1\right|,$ (7)
${\rm MAE}$ $=\frac{1}{M}\sum_{m=1}^{M}\frac{\left|N_{m}-\hat{N}_{m}\right|}{N_{m}},$ (8)
where $M$ is the total number of test videos, ${N}_{m}$ and $\hat{N}_{m}$ are the predicted and GT counts for the $m^{th}$ test video, respectively.
To better evaluate the performance of different models in recognizing actions with varying periods, we expand the evaluation metrics with three variants for the OBO and MAE metrics. We split the test video set into three categories based on the average single action period length: short-, medium-, and long-period test sets. We define videos with an average single action duration of fewer than 30 frames as belonging to the short-period test set, videos with an average action duration longer than 60 frames as the long-period test set, and the remaining videos as belonging to the medium-period test set. We compute the OBO and MAE metrics on each of these sets separately.
4.2 Implementation Details
We implement our approach with different backbone video feature extractors respectively, including TSN [41], I3D [4], and ViT [8]. For encoder-decoder Transformer, we set ${L}^{\text{enc}}=2$ and ${L}^{\text{dec}}=4$, both with 8-head attention mechanisms. The feature dimension is set to $C=512$. We set $Q=40$ queries in our method empirically. For more results with different feature channels $C$ and query numbers $Q$, please refer to the supplementary materials. The length of the video input is set to $T=512$ frames without down-sampling. We utilize the AdamW optimizer [27] with a learning rate of $0.002$, a batch size of $64$, and train the model for $80$ epochs. We set $\lambda_{\text{Hungarian}}=1.0$, $\lambda_{\text{L1}}=5.0$, $\lambda_{\text{gIoU}}=0.4$, $\lambda_{\text{ctrs}}=1.0$, confidence threshold $\alpha=0.2$. We provide further implementation details in the supplementary.
4.3 Comparison to State-of-the-arts
Table 2: Generalization comparison with SOTA TRC methods on UCFRep [47] dataset. MAE and OBO metrics for short-, medium-, and long-period actions are also reported.
Backbone ${\rm MAE}\downarrow$ ${\rm OBO}\uparrow$ ${\rm MAE_{s}}\downarrow$ ${\rm OBO_{s}}\uparrow$ ${\rm MAE_{m}}\downarrow$ ${\rm OBO_{m}}\uparrow$ ${\rm MAE_{l}}\downarrow$ ${\rm OBO_{l}}\uparrow$
RepNet [9] ResNet 0.5336 0.2984 0.6219 0.1739 0.4825 0.3600 0.4996 0.5000
TransRAC [14] ViT 0.6180 0.3143 0.6296 0.1951 0.5842 0.4250 0.6784 0.4118
Li et al. [19] 3D-ResNext 0.5227 0.3500 - - - - - -
Ours TSN 0.6016 0.2959 0.7069 0.0488 0.5777 0.4250 0.4039 0.5882
Ours I3D 0.5194 0.3980 0.4945 0.2195 0.5865 0.4250 0.4216 0.7647
Ours ViT 0.5435 0.4184 0.5657 0.1951 0.4625 0.5500 0.6804 0.6471
Table 3: Comparison of computational complexity. We benchmark all the methods with inference on the RepCountA dataset.
       $T$        Params (M)        FLOPs (G)
       RepNet [9]        64        20.47        92.40
       TransRAC [14]        64        14.43        100.19
       Ours        64        18.87        1.44
       Ours        512        18.87        6.06
Results on RepCountA dataset. We compare our proposed approach to the state-of-the-art methods on the RepCountA [14] dataset following previous work [14, 19] in Tab. 1. We compare with the SOTA action recognition [10, 26, 25], action segmentation [15] methods (top block), and TRC [9, 47, 14, 19] approaches (second block). We further adapt recent DETR-style action detection approaches [24, 45, 34] for the TRC task (third block). We change their output layers accordingly and train them on RepCountA.
As shown in Tab. 1, our approach significantly outperforms the state-of-the-art methods across actions of varying lengths. Specifically, the similarity-matrix-based methods [9, 14, 19] suffer from quadratic computation complexity. To manage this, they employ a sparse sampling strategy, ensuring reasonable content coverage within a limited temporal context window. However, this approach results in inferior performance for short, rapid action instances, as it tends to overlook the cycles. Conversely, the action detection approaches [24, 45, 34] are primarily developed for detecting action instances specific to particular classes. To adapt them to the TRC task, we modify their output layer to incorporate class-agnostic supervision accordingly. However, relying solely on class-agnostic supervision does not support them to dynamically identify repetitive cycles based on the input videos. As a result, their performance on long, slow action instances is generally inferior compared to dedicated TRC approaches. In contrast, our approach effectively balances the detection of actions at various speeds.
Results on UCFRep dataset. We also evaluate the generalization ability of our method. Following previous work [14, 19], we evaluate the model trained on the RepCountA dataset [14] on UCFRep [47] test set. Tab. 2 shows that our approach generally outperforms existing works, with more significant improvements observed in longer-period actions. The results demonstrate the effectiveness and generalization capability of our method.
Figure 3: Qualitative results on RepCountA dataset. Each colored block represents a GT or predicted action instance. TransRAC represents the results by density map, and the final count value is obtained by summing the values in the density map.
Figure 4: Visualization of failure case on RepCountA [14] dataset. Due to the excessive zooming in, the legs of the human body are truncated, making a large difference in the action motion feature, and resulting in several missed cycle counts. Each colored block represents a GT or predicted action instance.
Figure 5: Qualitative results on UCFRep dataset. Each colored block represents a GT or predicted action instance. TransRAC represents the results by density map, and the final count value is obtained by summing the values in the density map. The vertical lines in the labels represent the time points at which the actions begin since only the starting point annotations are provided in UCFRep [47].
Table 4: Effect of DAQ and ICL modules on RepCountA [14] dataset. (a) ablates DAQ strategy. (b) ablates the ICL strategy. (c) is our full model.
${\rm MAE}\downarrow$ ${\rm OBO}\uparrow$ ${\rm MAE_{s}}\downarrow$ ${\rm OBO_{s}}\uparrow$ ${\rm MAE_{m}}\downarrow$ ${\rm OBO_{m}}\uparrow$ ${\rm MAE_{l}}\downarrow$ ${\rm OBO_{l}}\uparrow$
(a) w/o DAQ 0.3542 0.4172 0.2624 0.2093 0.2515 0.5227 0.4864 0.4844
(b) w/o ICL 0.4035 0.4040 0.2448 0.2093 0.3106 0.4545 0.5740 0.5000
(c) Ours (full) 0.2809 0.4570 0.2411 0.1628 0.1792 0.5455 0.3776 0.5938
Efficiency. Additionally, we evaluate the computational complexity of our method in comparison to state-of-the-art methods. We benchmark all the methods with inference on the RepCountA dataset. For fairness of comparison, we calculate the Floating Point Operations (FLOPs) in the counting module for all the methods, excluding the video feature extraction process. We set the batch size to be $1$ and clip length to be $64$ for each method. Tab. 3 shows that our query-based method significantly reduces the computation load compared to the previous similarity-matric-based methods [9, 14], while also delivering superior accuracy. Importantly, the difference in computational demand is expected to widen with longer input clip lengths. This is because the computational complexity for similarity-matrix-based methods increases quadratically, whereas it grows linearly in our approach. Unfortunately, due to computational limitations, we are unable to obtain the results of RepNet [9] and TransRAC [14] when $T=512$.
4.4 Qualitative Results
We visualize the predictions of our approach and baseline method TransRAC [14] on the RepCountA [14] dataset in Fig. 3. TransRAC [14] represents action cycles using density maps, and the final count value is obtained by summing the values in the density map. However, this approach suffers from a lack of interpretability and finally results in miscounting. In contrast, our approach not only produces an accurate final count but also correctly localizes the action start and end positions (colored blocks) in most cases. In addition, our method exhibits robustness to changes in viewpoint, background noise, and sudden interruptions, e.g., in the second case of Fig. 3, we accurately estimate the time positions even with viewpoint changes as well as a sudden interruption in the middle of the timeline. Please refer to our supplementary video for more qualitative results. We present the video result of the first case in Fig. 3, which shows that while the subject performs pull-ups, there are moments of talking and arm waving. Our method selectively counts only the pull-up actions, as they are the primary focus and consistent with the main action instances of the video, excluding dissimilar actions like “waving”, validating the robustness and generalization ability of our method.
In addition, our query-based representation offers excellent interpretability, making it easy to identify the issues in case of failures. For example, Fig. 4 shows a typical failure case. Due to the excessive zooming in, the legs of the human body are truncated, making a large difference in the action motion feature, and resulting in several missed cycle counts.
We further illustrate the generalization performance of the proposed method on the unseen UCFRep [47] test set in Fig. 5. We directly apply our trained model and do not use UCFRep training data. Our model still accurately recognizes the action instances and gets the correct count in the challenging cases, indicating robust generalization ability. Specifically, the top case exhibits extreme viewpoint and lighting conditions, while the bottom case contains the action of soccer juggling which is not seen in the training set. We attribute the performance advantage to the proposed DAQ and ICL designs, which empower the model to adaptively adjust the action queries based on the input video features and effectively localize similar (repetitive) action instances, distinguishing them from the background noise actions.
4.5 Ablation Study
Effect of DAQ and ICL. We implement two ablated models to study the efficacy of the proposed DAQ and ICL designs. Tab. 4 presents the results on RepCountA [14] using the TSN backbone. In ablation (a), we substitute the proposed DAQ module with a static action query, where the action queries $\mathbf{Q}^{\text{act}}$ are also learnable variables, instead of using the action queries filtered out by the encoder and query selection modules. In ablation (b), we eliminate the ICL design among the action queries. The results demonstrate that both DAQ and ICL contribute to performance improvement, particularly in terms of counting medium and long actions.
Confidence threshold $\alpha$. We report the impact of different confidence thresholds for classifying “repetitive actions” on the final counts in Fig. 6. We illustrate the MAE (left) and OBO (right) curves for our method (green curve) and the TransRAC [14] approach (blue curve) across various confidence thresholds $\alpha$. The metrics for TransRAC are derived by binarizing its output density map and summing the results to obtain the final count. The results indicate that setting the threshold within the range of $0.2$ to $0.4$ yields similar performance for our method. Our method consistently outperforms TransRAC [14] (blue curve) by a considerable margin.
Figure 6: Results of different confidence thresholds of our method and TransRAC [14]. We depict the MAE (left) and OBO (right) curves of our method (green curve) and the TransRAC (blue curve) approach concerning different confidence thresholds. The metrics of TransRAC are obtained by binarizing the density map of TransRAC output and then summing to obtain the final count.
5 Conclusion
In conclusion, we provide an innovative perspective for tackling TRC tasks, which reduces computational complexity and maintains robustness across varying action lengths. To address open-set action categories, we propose DAQ for improved generalization, and ICL for recognizing repetitive actions and distinguishing them from distractions. Integrating DAQ and ICL, our method adaptively identifies contextually similar action instances. Experimental results on challenging benchmarks demonstrate our approach’s superiority over SOTAs in both accuracy and efficiency, while adeptly balancing diverse action speeds and video durations, establishing a solid foundation for practical implementations in real-world scenarios.
[…]
Acknowledgements This work was supported by National Science and Technology Major Project (2022ZD0114904).
[…]
Data Availability Statements. The RepCountA [14] and UCFRep [47] datasets that support the findings of this study are publicly available on GitHub at https://github.com/SvipRepetitionCounting/TransRAC and https://github.com/Xiaodomgdomg/Deep-Temporal-Repetition-Counting, respectively.